A ball is released from a window that is 10 meters above the ground. When
the ball leaves your hand, it is moving 8m/s at an angle 30
1
Expert's answer
2021-08-27T14:39:39-0400
"10=8\sin{30}t+0.5(9.8)t^2\\t=1.08\ s\\
x=8(1.08)\cos{30}=7.5\ m"